### v1

# X-ray bright AGN in local dwarf galaxies: insights from eROSITA

Abstract
Although supermassive black holes (SMBHs) reside in the heart of virtually every massive galaxy, it remains debated whether dwarf galaxies also commonly host SMBHs. Because low-mass galaxies may retain a memory of the assembly history of their black holes, probing the black hole occupation fraction of local dwarf galaxies might offer insights into the growth and seeding mechanisms of the first black holes. In this work, we exploit the Western half of the eROSITA all-sky survey (covering $20,000~{}\rm{deg^{2}}$) and compile a catalog of accreting SMBHs in local ($D<200$ Mpc) dwarf galaxies. After cleaning our sample from cosmic X-ray background sources, X-ray binaries, and ultraluminous X-ray sources, we identify 74 AGN-dwarf galaxy pairs. Using this large and uniform sample, we derive a luminosity function of dwarf galaxy AGN, fitting it with a power law function and obtaining ${\rm d}N/{\rm d}L_{\textup{X}}=(15.9\pm 2.2)\times L_{\textup{X}}^{-1.63\pm 0.05}$. Measuring the offset between the centroid of dwarf galaxies and the X-ray sources, we find that about $50\%$ of the AGN are likely off-nuclear, in agreement with theoretical predictions. We also compare the black hole-to-stellar mass relation of the AGN in our sample with the local and high-redshift relations, finding that our sources better adhere to the former. This suggests that local AGN across different mass scales underwent a similar growth history. Finally, we compare our sources with semi-analytical models: while our sample is too shallow to distinguish between different seeding models, it favors a growth mechanism linked to the star-formation rate of the host galaxy.
supermasive black holes – dwarf galaxies – accretion – X-ray active galactic nuclei
1 Introduction
Supermassive black holes (SMBHs) are among the most mysterious components of the universe. These SMBHs are found at the center of virtually every massive galaxy and have a profound influence on the evolution of their host galaxies from the earliest epochs to the present day (e.g. Croton et al., 2006; Hopkins et al., 2006). Deep optical surveys have revealed a population of optically bright quasars (i.e., SMBHs powered by accretion) at high redshifts ($z\sim 6$), indicating that black holes assembled rapidly, reaching masses of approximately $\sim 10^{9},{\rm M}_{\odot}$ within the first billion years after the Big Bang (e.g. Fan et al., 2006; Willott et al., 2007; Jiang et al., 2008; Mortlock et al., 2011; Venemans et al., 2013; Bañados et al., 2014). Recently, Chandra and JWST observations have detected black holes at even higher redshifts, when the universe was only 400-600 million years old (Larson et al., 2023; Bogdán et al., 2024; Kokorev et al., 2023; Maiolino et al., 2023; Kovács et al., 2024). Interestingly, these SMBHs are located in low-mass galaxies that exhibit similarities to present-day dwarf galaxies.
Figure 1: Aitoff projection of the source in the parent HECATE sample (in grey) and the matches with eRASS1 (red markers). The sources in our final sample are represented by filled red dots while the sources identified in the MILLIQUAS catalog and excluded from our sample are indicated by empty red circles.
The formation of SMBHs is a particularly intriguing question, with several promising pathways suggested: remnants of Population III stars (Bromm & Loeb, 2003); mergers in nuclear star clusters (Portegies Zwart et al., 2004; Gürkan et al., 2004; Natarajan, 2021); direct collapse of pre-galactic gas disks where fragmentation and star formation are suppressed (Lodato & Natarajan, 2006; Begelman et al., 2006); or dark stars powered by dark matter annihilation (Spolyar et al., 2008). There are two promising avenues to distinguish between various theoretical models. One approach is to observe SMBHs in the early universe ($z\gtrsim 9$), close to their seeding epoch. Alternatively, we can investigate the occupation fraction and properties of SMBHs in dwarf galaxies at the present epoch.
Studying the population of SMBHs in nearby dwarf galaxies offers a powerful and complementary method to probe the seeding mechanisms and growth history of the first black holes. While SMBHs in today’s massive galaxies ($>10^{10}\ \rm{M_{\odot}}$) did not retain a memory of their origins due to the intense merger activity of their host galaxies, the black hole occupation fraction of local dwarf galaxies ($10^{8.5-9.5}\ \rm{M_{\odot}}$) is sensitive to the black hole seeding mechanism (Volonteri & Begelman, 2010). If SMBHs formed through the collapse of Population III stars, every dwarf galaxy in the local universe is expected to host an SMBH. However, if black holes formed through the direct collapse of massive gas clouds, only about half of present-day dwarf galaxies are predicted to harbor an SMBH (see e.g. Greene 2012).
Over the past years, various methods have been employed to detect accreting SMBHs (or Active Galactic Nuclei – AGN) in dwarf galaxies (see Mezcua 2017; Greene et al. 2020; Reines 2022 and references therein). The most common approach for systematically identifying AGN in dwarf galaxies is based on optical spectroscopy, utilizing emission-line diagnostics to distinguish between star formation-related and accretion-powered (i.e., AGN) activity (Greene & Ho, 2004, 2007; Reines et al., 2013; Moran et al., 2014; Baldassare et al., 2015; Chilingarian et al., 2018). While highly efficient, identifying $\gtrsim 600$ AGN in dwarf galaxies to date, this technique is biased towards objects accreting at a high Eddington ratio and those hosted in solar and super-solar metallicity galaxies (e.g., Greene et al. 2020; but see Polimera et al. 2022 and Mezcua & Domínguez Sánchez 2024 for ways to correct for these effects).
Alternative methods to identify SMBHs in dwarf galaxies rely on optical, UV, and infrared variability (Baldassare et al., 2018, 2020; Martínez-Palomera et al., 2020; Ward et al., 2022; Arcodia et al., 2024), coronal lines (Satyapal et al., 2007; Cann et al., 2018, 2021; Salehirad et al., 2022; Reefe et al., 2023), radio emission (Mezcua et al., 2018; Reines et al., 2020; Davis et al., 2022), and X-ray observations (Schramm et al., 2013; Lemons et al., 2015; Pardo et al., 2016; Mezcua et al., 2018; Birchall et al., 2020; Latimer et al., 2021; Bykov et al., 2024). These last two methods are particularly effective in identifying AGN due to the high nuclear-to-host contrast that these bands offer: high angular resolution radio observations allow the separation of the AGN from other radio-emitting components of the galaxy, and the X-ray spectrum of AGN is distinct from other stellar X-ray sources (Ho, 2008; Merloni, 2016). However, a sensitive volume-limited radio survey of dwarf galaxies in the nearby universe is prohibitively expensive (Reines et al., 2020). Similarly, X-ray observational campaigns have, to date, been either too shallow and with poor angular resolution, or too limited in sky coverage.
The game-changer in this domain is eROSITA (extended ROentgen Survey with an Imaging Telescope Array, Predehl et al. 2021), the main instrument aboard the German-Russian satellite Spektrum Roentgen Gamma (SRG) (Sunyaev et al., 2021). eROSITA provides today’s most sensitive all-sky survey in the X-ray band, and eRASS1, the first data release, comprises more than 930,000 individual sources across the Western half of the sky (Merloni et al., 2024). In this work, we leverage the opportunity offered by eRASS1 to build a sample of uniformly selected AGN in local dwarf galaxies, thereby probing their black hole occupation fraction.
A similar effort has been conducted by Bykov et al. (2024), who utilized eROSITA data from the Eastern half of the sky and the MPA-JHU galaxy catalog, based on the SDSS DR12 (Alam et al., 2015), with promising results: they obtained a sample of 99 X-ray source-dwarf galaxy pairs, of which 82 were identified as AGN.
The paper is organized as follows: in Section 2 we describe the sample-building procedure and cleaning criteria; in Section 3 we analyze the properties of the sources in our sample; in Section 4 we compare our sample with semi-analytical predictions and in Section 5 we draw our conclusions.
2 Sample selection and properties
2.1 Defining the galaxy sample
To search for AGN candidates located in dwarf galaxies within the eRASS1 footprint, we cross-matched the positions of approximately 930,000 X-ray sources in the eRASS1 catalog (Merloni et al., 2024) with those of local dwarf galaxies. To define the sample of nearby dwarf galaxies, we relied on the HECATE catalog (Kovlakas et al., 2021), an all-sky, value-added catalog that includes information about the size, distance, and star-formation rate of more than 200,000 local galaxies. It is more than $50\%$ complete in terms of B-magnitude up to $\approx 170$ Mpc. Filtering the HECATE catalog for local dwarf galaxies (within $D<200$ Mpc and with stellar masses of $10^{6}~{}{\rm M_{\odot}}<M_{\rm gal}<3\times 10^{9}~{}{\rm M_{\odot}}$) resulted in a sample of 5,775 objects. We adopted a radius of $15\arcsec$ to cross-match the positions of eRASS-detected X-ray sources with the dwarf galaxies. This radius corresponds to $\approx 10.5$ kpc at the mean distance of the galaxy sample, which is approximately three times the mean radius of the dwarf galaxies. We thus obtained 120 X-ray source-dwarf galaxy pairs. The location of the sources in the parent HECATE sample and the results of our cross-match are shown in Figure 1.
2.2 Chance alignment with background AGN
A fraction of these matches likely result from chance alignments between dwarf galaxies and background AGN: cosmic X-ray background sources could be projected at the positions of the dwarf galaxies. We used two approaches to determine the number of such sources. First, we carried out 5,000 Monte Carlo simulations by generating 5,775 random coordinates representing the dwarf galaxy population in these simulations. Then we cross-matched the generated coordinates with the eRASS1 catalogue adopting the same match radius of $15\arcsec$. On average, we found $14\pm 4$ matches which can be attributed to chance alignment. The distribution of the matches obtained through the Monte Carlo procedure is shown in Figure 2.
Second, we compared the locations of the X-ray source-galaxy pairs to the MILLIQUAS catalog (Flesch, 2023), which includes nearly one million quasars. This comparison yielded 17 matches, aligning well with the results of the Monte Carlo simulations. We then removed these 17 sources from our sample, reducing it to 103 X-ray source-dwarf galaxy pairs.
Figure 2: Histogram of the number of matches obtained in 5,000 iterations of randomly generated coordinated and the eRASS1 catalog, the vertical red lines indicate the average number of matches (solid) and the $1\sigma$ uncertainty interval (dashed).
2.3 Unresolved X-ray binaries
The population of X-ray binaries can significantly contribute to the X-ray emission of galaxies. Due to the broad PSF of eROSITA and the shallow depth of the eRASS1 survey, distinguishing these X-ray binaries from AGN emission is challenging. Because the collective emission from X-ray binaries correlates with the galaxy’s stellar mass (for low-mass X-ray binaries) and star-formation rate (for high-mass X-ray binaries), we probed whether the population of X-ray binaries could account for the X-ray emission observed in the studied dwarf galaxies. To this end, we used the galaxy-wide X-ray binary emission scaling relation (Lehmer et al., 2010):
$L_{2-10\,{\rm keV}}\,{\rm(erg/s)}=\alpha M_{\textup{gal}}+\beta{\rm SFR}\ ,$
where $\alpha=(9.05\pm 0.37)\times 10^{28}\,{\rm erg/s/M_{\odot}}$ and $\beta=(1.62\pm 0.22)\times 10^{39}\,{\rm erg/s/(M_{\odot}/y)}$. This relation has a dispersion of 0.34 dex.
The HECATE catalog provides stellar masses for all galaxies in our sample and SFR values for 47 of them. For the remaining sources, we computed their SFR using the relation suggested by Rieke et al. (2009): ${\rm SFR}=7.8\times 10^{-10}\,L_{24\,\mu}/\rm{L}_{\odot}{\rm(erg/s)}$. For 57 sources, we used the $22,\mu$ luminosity (W4 Allwise band) instead of the $24,\mu$ luminosity, and for two galaxies, we used the $25,\mu$ luminosity (IRAS 25), assuming that $L_{24,\mu}/L_{25,\mu}\approx L_{24,\mu}/L_{22,\mu}\approx 1$. Finally, for two galaxies without infrared photometry, we assumed the average SFR of our parent sample.
The formula provided by Lehmer et al. (2010) derives the luminosity in the $2-10$ keV band, so we converted the $0.2-2.3$ keV luminosity provided by the eRASS1 catalog to the $2-10$ keV band assuming a Galactic absorption of $3\times 10^{20}$ cm-2 and a power law with photon index $\Gamma=1.8$.
For 15 galaxies, the X-ray emission from the eRASS1 sources is comparable to the X-ray emission expected from the population of unresolved X-ray binaries. For the remaining galaxies, the X-ray sources are more than 2.2 times brighter than expected from the population of X-ray binaries, making it unlikely that the eRASS-detected X-ray sources originate from X-ray binaries.
Figure 3: Observed vs expected X-ray luminosity from X-ray binaries for our sample of dwarf galaxies. The dashed line indicates the bisectrix. The cumulative emission from X-ray binaries may account for the observed eROSITA X-ray emission in the sources highlighted in red. These galaxies were removed from our sample.
2.4 Resolved ultra-luminous X-ray sources
The sources in the high-luminosity tail of the distribution of low-mass and high-mass X-ray binaries (LMXBs/HMXBs), appearing as single, resolved sources, might still contaminate our sample of X-ray source-dwarf galaxy pairs. These X-ray sources are expected to dominate the low-luminosity tail of our sample. To be conservative, we removed all sources from our sample with luminosity $L_{\rm X}<10^{39}\ \rm{erg\ s^{-1}}$ in our reference $0.2-2.3$ keV band. Assuming all these were HMXBs, we then inferred the number of residual contaminants in the rest of the sample.
To estimate this number, we adopted the HMXB luminosity functions by Mineo et al. (2012) and Geda et al. (2024). The former better samples the high-luminosity end of the distribution, as it is based on a larger sample of sources, but it is not calibrated for dwarf galaxies. Conversely, the latter is calibrated for dwarf galaxies but is heavily affected by stochasticity at its high-luminosity end. Figure 4 shows the luminosity distribution of the X-ray sources in our sample, with the high-mass X-ray binary luminosity functions superimposed, normalized for the number of sources with $L_{\rm X}<10^{39}\ \rm{erg\ s^{-1}}$. Based on these models, we expect that our final sample of 78 sources will be contaminated by 24242–42 – 4 HMXBs with a luminosity exceeding $10^{39}\ \rm{erg\ s^{-1}}$. Even accounting for an incompleteness fraction of about $50\%$ (see Section 3.3 for further details), the number of expected contaminants is $\lesssim 6$.
Figure 4: Distribution of the observed X-ray luminosity of the AGN in our sample. Dashed and dotted lines indicate the predicted HMXBs populations normalized to the number of sources with a luminosity $L_{\rm X}<10^{39}\ \rm{erg\ s^{-1}}$.
Figure 5: A representative subset of our sample of AGN hosted in local dwarf galaxies. The three-color composite images were obtained using the SDSS gri-filters. The X-ray counts are shown with the white (two counts) and red (one count) pixels. The cross and the small magenta circle indicate the centroid and positional uncertainty ($\approx 5^{\prime\prime}$) of the X-ray source taken from the eRASS1 point source catalog (Merloni et al., 2024), respectively. The large cyan circle shows the PSF size ($r=26^{\prime\prime}$) of eROSITA in scanning mode.
2.5 Crossmatch with archival X-ray data
A subset of the sources in our sample was previously observed with XMM-Newton and/or Chandra, and data are available in the respective archives [^1] https://nxsa.esac.esa.int/nxsa-web/#search, https://cda.harvard.edu/chaser/mainEntry. One source represents a spurious detection, not associated with a point-like source but rather with the diffuse emission of a nearby galaxy cluster, and was hence removed from our sample. Four sources in our sample are identified as ultraluminous X-ray sources (ULXs), i.e. HMXBs with a luminosity exceeding $10^{39}\ \rm{erg\ s^{-1}}$, based on their multi-wavelength emission.
Three sources were identified as candidate AGN due to their infrared and X-ray emission, and their nature has been further investigated with Chandra pointed observations. In two cases, the Chandra observations revealed compact sources of hard X-ray photons in the nuclear regions of the host galaxies, confirming the AGN nature of their emission (Dudik et al., 2005; Latimer et al., 2021). In one case, however, the compact source lies outside the nuclear region of its host galaxy. While this might suggest that the source is a ULX (Thygesen et al., 2023), it is also possible that it is an off-nuclear AGN, as a fraction of dwarf galaxies ($\sim 50\%$) are expected to host wandering SMBHs outside the nuclear region (Reines et al., 2020; Bellovary et al., 2021).
Finally, five additional sources fall within the footprints of either XMM-Newton or Chandra observations. While detected as point-like sources, they are too far off-axis for their locations to be precisely assessed within their host galaxies, although their X-ray spectra are consistent with being AGN. A detailed description of all the sources mentioned above is presented in Appendix B.
2.6 Final sample of dwarf galaxy-X-ray source pairs
Our final sample, after removing the four known ULXs, is composed of 74 dwarf galaxies with an eROSITA-detected source. In Figure 6 we show the X-ray luminosity of the sources as a function of the redshift of the galaxies. Postcard images of a representative set of three of these candidates are shown in Figure 5. All the relevant properties of each dwarf galaxy-X-ray source pair are reported in Table 2.
3 Results
3.1 X-ray hardness ratios
To further characterize the properties of the eROSITA-detected X-ray sources, we first probe their spectral properties. Because the bulk of the sources in our sample only have a handful of X-ray counts, we cannot obtain reliable hardness ratios for most of the individual sources. Therefore, we compute the sample-average X-ray hardness ratio. To this end, we define the hardness ratio as $\rm{HR}=S/M$ where the number of counts is measured in the soft (S=$0.5-1$ keV) and medium bands (M=$1-2$ keV), respectively. We obtained a hardness ratio of $S/M=0.98_{-0.06}^{+0.04}$, where the uncertainties were computed using a Bayesian approach with the BEHR code (Park et al., 2006). Considering the eROSITA effective area curve, this measured value is compatible with a power-law model with a slope of $\Gamma=1.8$ seen through a column density of $N_{\textup{ H}}=3\times 10^{20}$ cm-2. This spectral model is in good agreement with the typical slope of AGN detected in the eROSITA early data release (eFEDS, Liu et al. 2022) and disfavor the scenario in which the X-rays we detect come from diffuse emission.
Figure 6: X-ray luminosity in the $0.2-2.3$ keV eRASS1 band as a function of the redshift of each source in our final sample of 74 AGN.
3.2 Off-nuclear SMBHs
Another key piece of information that helps us better characterize our sample and compare our results with theoretical predictions is the offset between the X-ray source and the host galaxy centroid. As argued for one of the sources with a Chandra follow-up observation, the most natural explanation for an off-nuclear X-ray bright source is a ULX scenario (see e.g. Kaaret et al. 2017; Fabrika et al. 2021 for reviews on the topic). However, not only do we detect sources in excess of the ones expected from HMXB/ULX luminosity functions, but given the shallow gravitational potential of dwarf galaxies, the fraction of off-nuclear AGN is expected to be as large as one-third (Reines et al., 2020; Bellovary et al., 2021).
Figure 7 shows the offset of the X-ray source position with respect to its host galaxy centroid. The bulk of our sample has a small offset, with only about 10 sources at a distance larger than $2$ kpc from their host galaxy centroid. Once the positional error of the X-ray sources (POS_ERR from the eRASS1 catalog) is taken into account, the location of the X-ray source is compatible with the nucleus of its host galaxy for about $50\%$ of the objects in our sample, and for more than $80\%$, it is compatible within two times the positional error.
These results align with theoretical predictions and observational evidence (Bellovary et al., 2019; Reines et al., 2020; Bellovary et al., 2021) for AGN in dwarf galaxies (as also shown in blue in Figure 7), although some caveats must be considered. The large PSF of eROSITA in scanning mode ($\approx 26\arcsec$) prevents a precise location of the X-ray sources, and dwarf galaxies often exhibit irregular morphology, making it sometimes difficult to identify their centers (Pasetto et al., 2003; Lazar et al., 2024). To conclusively determine the location of the X-ray sources, follow-up high spatial resolution observations with Chandra would be needed.
Figure 7: Offset of the X-ray source position with respect to the host galaxy centroid in kpc. The red dashed line indicates the average radius of the dwarf galaxies in our sample. The blue histogram shows the prediction by Bellovary et al. (2019), rescaled to match our sample size.
3.3 X-ray luminosity function
To further characterize the sources, infer properties of the bulk population of AGN in dwarf galaxies, and compare them with populations of other X-ray emitting sources, we constructed the X-ray luminosity function of our sample. To correct for the eROSITA incompleteness, we weighted each luminosity bin by the probability of detecting an X-ray source in that bin. The probability of detecting a source with luminosity $L_{\textup{X}}$ was computed by simulating $N$ sources distributed in redshift like the sources in our parent sample. Using the provided sensitivity map, we calculated how many sources, $D$, eROSITA would detect. The probability of detecting a source with luminosity $L_{\textup{X}}$ is simply $D/N$.
The X-ray luminosity function we obtained is shown in Figure 8. We also show the luminosity functions of HMXBs/ULXs from previous works (Mineo et al., 2012; Geda et al., 2024). A direct comparison between these and the luminosity function derived from our sample clearly shows that X-ray source-dwarf galaxy pairs in our sample represent a population of sources with significantly different behavior in terms of X-ray emission. We fit our luminosity function with a power-law model and obtained the following:
${\rm d}N/{\rm d}L_{\textup{X}}=(15.9\pm 2.2)\times L_{\textup{X}}^{-1.63\pm 0.05},$ (1)
where $L_{\textup{X}}$ is expressed in units of $10^{41}\ \rm{erg\ s^{-1}}$. The luminosity function we obtained, although slightly steeper, is compatible with those reported by Birchall et al. (2020) (based on XMM-Newton X-ray data) and Bykov et al. (2024) (based on the eROSITA data of the Eastern sky), both of which used SDSS data as dwarf galaxies parent sample. The difference might stem from the slightly different selection criteria and the luminosity range we considered: both authors report a flattening of the luminosity function below $10^{39}\ \rm{erg\ s^{-1}}$, which we are not sensitive to as we excluded all sources below that threshold.
Figure 8: X-ray luminosity function of our sample of AGN in dwarf galaxies (black data points), the raw luminosity function, before the completeness correction is also shown (grey data points). The black solid line and grey shaded region indicate our best fit and its dispersion. The blue and red dashed lines indicate the luminosity functions of HMXBs/ULXs (Mineo et al., 2012; Geda et al., 2024).
3.4 Galaxy-to-black hole mass relation
To place our results in a broader context, we compare the black hole mass ($M_{\bullet}$) of the AGN in our sample, inferred from their X-ray luminosity, to the mass of their host galaxies ($M_{\star}$). The recent detection of AGN at high redshift, powered by SMBHs with a mass equal to $\approx 10-100\%$ of their host galaxy (Pacucci et al., 2023; Bogdán et al., 2024; Maiolino et al., 2023), has prompted theoretical efforts to reconcile these observations with the local $M_{\bullet}-M_{\star}$ relation. This local relation typically accounts for SMBHs with a mass $\approx 0.1\%$ of the galaxies they inhabit (Reines & Volonteri, 2015). It might be possible to reconcile this discrepancy by invoking observational biases, such as the difficulty in detecting high-redshift, low-mass (and hence low-luminosity) AGN (Li et al., 2024).
Exploiting the analogies between high-redshift AGN and AGN hosted in low-mass galaxies, we further investigate these effects by studying the $M_{\bullet}-M_{\star}$ relation in our dwarf galaxy sample. We infer the SMBH mass from the X-ray luminosity of the AGN. To this end, we employ the bolometric correction $L_{\textup{bol}}/L_{\textup{2-10keV}}=16.7$ (see e.g., Lusso et al. 2012; Duras et al. 2020) and assume that all black holes in our sample accrete at $1\%$ of their Eddington ratio. The inferred black hole masses are in the range of $\sim 10^{4}-10^{1}0^{7}\ \rm{M_{\odot}}$. Figure 9 shows the $M_{\bullet}-M_{\star}$ relation for the AGN in our sample. The blue line and shaded region indicate the local relation for AGN and its dispersion (Reines & Volonteri, 2015). The red and green lines and shaded regions show the high-redshift relations proposed by Pacucci et al. (2023) and Li et al. (2024), respectively.
Our sample of AGN in dwarf galaxies is compatible with the low-redshift $M_{\bullet}-M_{\star}$ relation, although it shows significant scatter and is biased towards higher luminosities at the low-mass end of the plot (the dotted line indicates the lower limit on the detectable black hole mass). This suggests that local AGN hosted in dwarf galaxies share the same $M_{\bullet}-M_{\star}$ relation as more massive $z=0$ galaxies, in agreement with previous results (Schutte et al., 2019).
Figure 9: $M_{\bullet}-M_{\star}$ for the AGN in our sample. The blue and red lines indicate the low- and high-redshift relations, respectively, with the shaded regions indicating their dispersion. The dotted lines indicate the limit of the parameter space we explored in our analysis: the lower limit on the detectable BH mass, corresponding to a luminosity limit of eRASS1 at the average distance of our sample and the upper limit of the galaxy mass. The magenta marker indicates UHZ1, a $z=10.1$ galaxy with a Chandra-detected accreting over-massive black hole (Bogdán et al., 2024).
4 Comparison with semi-analytical models
Semi-analytical models (SAM) represent a powerful tool for studying BH seeding mechanisms as well as for exploring growth modes of BHs (Somerville et al., 2008; Bower et al., 2010; Ricarte & Natarajan, 2018a, b). In this work, we compare the observed BH occupation fraction with mock catalogs of AGN built following the approach described in Chadayammuri et al. (2023) based on the SAMs developed by Ricarte & Natarajan (2018b).
These SAMs consider two seeding mechanisms and three different mass growth modes, combined to build six mock catalogs. The seeding mechanisms comprise either rare “heavy” seeds, typically resulting from direct collapse (DC) scenarios, distributed following Lodato & Natarajan (2006) and Volonteri & Begelman (2010), or abundant “light” seeds with a distribution motivated by cosmological simulations (Hirano et al., 2015). This second scenario is motivated by cosmological simulations where BHs are produced by the death of population III (popIII) stars and the mass distribution of the seeds is represented by a power law spanning from 30 to 100 solar masses (Hirano et al., 2015).
Each seeding mechanism is coupled with one of three mass growth modes, detailed below.
  • Power Law (PL): When a major merger occurs, central BHs grow at the Eddington rate until they reach the $M_{\bullet}-\sigma$ relation. Otherwise, they accrete according to a universal Eddington ratio distribution function.
  • AGN Main Sequence (AGNms): When a major merger occurs, central BHs grow at the Eddington rate until they reach the $M_{\bullet}-\sigma$ relation. Otherwise, they accrete at $10^{-3}$ times the star formation rate.
  • Broad Line Quasar (BLQ): When a major merger occurs, central BHs grow at Eddington ratios drawn from observed broad line quasars, allowing occasional super-Eddington accretion. Otherwise, the accretion rate is zero. [^2] This was shown to be adequate for reproducing quasar luminosity functions (Ricarte et al., 2019), but fails to produce enough faint AGN (Chadayammuri et al., 2023).
We can probe growth modes and seeding mechanisms by comparing our observed AGN properties with the mock catalogs built upon the described SAMs. Following Chadayammuri et al. (2023), for each SAM, we build the corresponding mock catalog by computing the expected X-ray luminosity based on the mass and mass accretion rate of each BH. We assume an efficiency in converting mass accretion rate into luminosity of $10\%$ and a bolometric-to-X-ray luminosity correction of $10\%$, which are both common and adequate choices (see e.g., Lusso et al. 2012; Duras et al. 2020). We then filter the obtained mock catalogs using the same luminosity, mass, and distance criteria adopted in the construction of our AGN sample in dwarf galaxies. This approach allows for a direct comparison between the filtered mock catalogs and our observed sample.
Figures 10 and 11 present our main results, showing the normalized distributions of host-galaxy masses and luminosities for our sample (in grey) and the six mock catalogs built on different models (colored dashed lines). Three models roughly reproduce the distribution of observed host-galaxy mass: DC+AGNms, popIII+AGNms, and popIII+BLQ. However, the popIII+BLQ model fails to predict the observed distribution of AGN luminosities. Even the two best-performing models (DC+AGNms and popIII+AGNms) fail to reproduce the observed luminosity distribution. They predict distributions heavily skewed toward high luminosities, peaking at about $10^{41.5}\ \rm{erg\ s^{-1}}$, while the observed distribution peaks roughly at $10^{39.5}\ \rm{erg\ s^{-1}}$. This suggests that the preferred accretion mode is the one adopted in the AGNms model, but it also highlights the need for more theoretical and modeling efforts to accurately reproduce the observed AGN luminosity distribution.
Figure 10: Normalized distribution of the AGN host galaxies masses for our sample (in grey) and the six mock catalogs obtained from the described SAM (dashed colored lines).
These results offer intriguing insights into the physics of SMBHs and their growth history. Dwarf galaxies typically experience fewer major mergers in their recent assembly compared to their more massive counterparts, making them more sensitive to the “steady mode” of accretion. Our results suggests that a good model for this steady mode could involve accretion tied to the star-formation rate of the host galaxy.
The results by Chadayammuri et al. (2023) were obtained using the AMUSE survey (Gallo et al., 2008; Miller et al., 2015), a volume-limited survey aimed at identifying AGN in early-type galaxies. Chadayammuri et al. (2023) concluded that data from the AMUSE survey favor the “heavy-seed” scenario but are too shallow to completely rule out the “light seed” scenario. This conclusion is further confirmed by our analysis. While we can effectively discriminate between accretion modes, our X-ray-selected data sample is too limited to probe different BH seeding mechanisms. This picture will not change significantly with the release of additional eROSITA scans: although the average detection sensitivity will improve by a factor of $\approx 2.5$, the low-luminosity tail of the AGN population will remain inaccessible. In this luminosity regime, LMXBs, HMBXs and ULXs will dominate, and the poor angular resolution of eROSITA does not allow for their unambiguous identification.
Figure 11: Normalized distribution of the AGN X-ray luminosity of our sample (in grey) and the six mock catalogs obtained from the described SAM (dashed colored lines).
5 Conclusions
Harnessing the recent release of eROSITA data of the Western part of the sky (which covers 20,000 $\rm{deg^{2}}$, we compiled a large and uniform sample of X-ray-detected AGN in local dwarf galaxies. After excluding all possible contaminants, our final sample consists of 74 X-ray source-dwarf galaxy pairs.
We characterized the sources in our sample by investigating their X-ray hardness ratio, their location within the host galaxy, luminosity function, and galaxy-to-BH mass relation. The average X-ray hardness ratio of our sample is compatible with a power-law model with a slope of $\Gamma=1.8$, as expected from a population of AGN. While the broad PSF of eROSITA prevents us from precisely locating the X-ray sources within their host galaxies, we found that a substantial fraction ($\approx 50\%$) of the sources are likely off-nuclear, in line with theoretical predictions. We also constructed a precise X-ray luminosity function for AGN in dwarf galaxies, which we fit with a power-law function. Finally, we compared the BH mass of our sources (inferred from their X-ray luminosity) with their host galaxy stellar mass. We found that the ratio of BH-to-galaxy mass for our sample is in good agreement with the local relation for AGN hosted in more massive galaxies.
To place our results in context, we compared the mass and luminosity distributions of our sample of AGN in dwarf galaxies with predictions from semi-analytical models. Based on this analysis we concluded that (a) the preferred mode of accretion appears to be one in which the accretion rate is linked to the star-formation rate of the host galaxy; (b) no model can convincingly reproduce the observed distribution of AGN luminosities, suggesting that more modeling effort is needed to better describe the behavior of AGN in dwarf galaxies; and (c) our sample is too shallow to distinguish between different seeding mechanisms, highlighting the necessity of deeper X-ray surveys to convincingly determine the origin of the observed SMBHs.
Á.B. acknowledges support from the Smithsonian Institution and the Chandra Project through NASA contract NAS8-03060. This work was supported in part by the Black Hole Initiative at Harvard University, made possible through the support of grants from the Gordon and Betty Moore Foundation and the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of the Moore or Templeton Foundations.
Appendix A Properties of the selected sources
We report in Table 2 the properties of the 74 AGN in dwarf galaxies we selected.
Table 1: Properties of the 74 AGN selected.
[…]
PGC135069 52.2929 -22.0213 52.2929 -22.0216 0.97 24.5 9.41 $6.55\pm 0.47$ 0
WINGSJ043139.17-612330.1 67.9132 -61.3916 67.9166 -61.3901 1.87 41.03 8.22 $11.37\pm 0.67$ 0
PGC200277 179.382 32.3408 179.3816 32.3399 3.98 48.36 9.02 $3.1\pm 1.09$ 1
6dFJ1551458-091534 237.9408 -9.2594 237.9406 -9.2597 2.01 75.19 8.87 $24.03\pm 3.99$ 1
ESO411-012 11.7126 -31.541 11.7127 -31.5383 4.89 19.71 7.66 $0.22\pm 0.09$ 0
PGC2822735 25.2132 -32.7364 25.2124 -32.7355 3.96 166.27 9.45 $64.99\pm 12.69$ 0
6dFJ0142532-432355 25.7217 -43.3986 25.7214 -43.3983 1.83 19.22 7.78 $1.72\pm 0.21$ 1
NGC1034 39.5584 -15.8091 39.5571 -15.8082 5.05 19.28 9.15 $0.27\pm 0.09$ 0
6dFJ0245228-240617 41.3451 -24.1046 41.3445 -24.1048 1.39 101.61 9.24 $72.89\pm 7.01$ 0
ESO479-025 40.533 -24.1292 40.533 -24.1275 5.09 16.9 9.31 $0.19\pm 0.07$ 0
PGC3213080 42.5881 -29.5708 42.5866 -29.571 6.68 76.4 8.76 $2.57\pm 1.06$ 1
PGC012154 49.0887 -25.8548 49.0894 -25.8546 5.3 19.66 9.37 $0.25\pm 0.08$ 1
PGC693568 47.0262 -32.3275 47.0256 -32.327 3.28 65.84 8.75 $5.14\pm 1.12$ 1
PGC540963 50.5536 -44.0946 50.5543 -44.0946 5.37 122.51 9.28 $3.66\pm 1.71$ 1
PGC134076 53.7236 -24.7909 53.7241 -24.7917 4.62 21.76 8.45 $0.19\pm 0.08$ 1
6dFJ0342278-260243 55.6158 -26.0452 55.6171 -26.0459 11.14 21.4 8.23 $0.18\pm 0.08$ 1
PGC3315038 63.0987 -54.4213 63.0968 -54.4219 6.25 173.09 9.47 $4.92\pm 2.21$ 0
6dFJ0421194-313805 65.3308 -31.6346 65.3303 -31.6341 1.28 32.95 8.67 $8.19\pm 0.78$ 0
PGC126927 66.4561 -61.1529 66.4528 -61.1535 5.46 68.36 8.88 $0.75\pm 0.31$ 0
ESO118-034 70.0728 -58.7443 70.0717 -58.7467 3.37 13.65 9.27 $0.21\pm 0.03$ 0
PGC928553 73.1952 -14.1312 73.1948 -14.1297 3.67 39.15 9.37 $0.95\pm 0.35$ 0
PGC016675 76.1302 -16.5847 76.1323 -16.5854 5.16 43.61 9.37 $0.59\pm 0.31$ 0
PGC016986 78.8377 -26.468 78.8379 -26.4709 3.8 50.47 8.71 $1.38\pm 0.52$ 0
UGC03282 79.4064 6.7994 79.4075 6.8012 5.59 80.21 9.42 $3.47\pm 1.64$ 0
6dFJ0532495-322925 83.2062 -32.4904 83.2047 -32.4939 4.64 148.13 9.47 $10.68\pm 4.07$ 0
PGC917425 84.0581 -14.977 84.0584 -14.9759 4.91 26.12 8.82 $0.62\pm 0.2$ 1
2MASXJ05403594-5436417 85.1497 -54.6116 85.1542 -54.6139 3.98 32.35 9.43 $0.38\pm 0.11$ 0
PGC148318 90.956 -33.1912 90.9581 -33.1915 5.65 36.89 9.17 $0.54\pm 0.36$ 0
PGC075555 93.0592 -38.7731 93.0568 -38.7732 4.78 21.07 8.43 $0.2\pm 0.09$ 0
PGC483594 92.438 -48.7986 92.4384 -48.8023 4.04 122.17 9.27 $10.04\pm 2.95$ 0
PGC019486 100.792 -76.5594 100.7784 -76.5586 5.15 50.73 9.13 $0.43\pm 0.16$ 0
PGC088580 97.7514 -56.6482 97.7462 -56.6468 10.4 142.62 9.41 $25.28\pm 6.02$ 0
PGC747791 102.7671 -27.9423 102.7673 -27.9426 4.24 31.13 9.1 $0.93\pm 0.32$ 1
PGC2054090 120.4398 34.7579 120.4402 34.7612 6.77 65.89 9.31 $2.13\pm 1.27$ 0
PGC023484 125.6199 -1.1028 125.6198 -1.1023 1.9 62.38 9.16 $34.09\pm 4.47$ 1
PGC3093997 130.9451 -17.6794 130.944 -17.6782 6.28 22.81 9.34 $0.43\pm 0.2$ 1
PGC024469 130.7008 14.2652 130.6996 14.2665 7.11 29.99 9.45 $0.44\pm 0.26$ 1
PGC153352 134.7511 -18.3842 134.7513 -18.3851 4.46 26.86 9.25 $0.6\pm 0.23$ 1
NGC2777 137.6744 7.2067 137.6738 7.2045 4.51 25.7 9.33 $0.74\pm 0.27$ 0
IC0549 145.1801 3.9592 145.1792 3.9598 6.69 22.7 8.93 $0.45\pm 0.2$ 1
PGC2077238 145.8306 36.2369 145.8302 36.2392 5.73 87.52 9.46 $4.91\pm 2.53$ 0
AGC193818 147.9617 13.9109 147.961 13.9144 14.23 98.06 9.39 $31.27\pm 9.12$ 1
NGC3125 151.6388 -29.9353 $1$29.9371 3.14 14.98 9.37 $0.49\pm 0.12$ 0
Table 2: Continued.
[…]
6dFJ1024202-201458 156.0839 -20.2494 156.0842 -20.2494 3.34 69.73 8.96 $15.63\pm 3.48$ 1
IC2604 162.3545 32.7728 162.354 32.7749 5.56 25.1 8.97 $0.44\pm 0.2$ 0
UGC05989 163.1327 19.7923 163.1322 19.7929 3.89 16.07 8.78 $0.51\pm 0.14$ 1
PGC096512 173.696 -59.2293 173.6943 -59.2298 5.91 18.51 8.48 $0.13\pm 0.05$ 0
NGC4016 179.6209 27.5288 179.6184 27.53 6.43 50.62 9.46 $1.6\pm 0.86$ 0
ESO321-018 183.9762 -38.0934 183.9723 -38.0909 4.25 41.79 9.14 $0.93\pm 0.37$ 0
ESO267-032 183.6912 -43.5649 183.6885 -43.565 5.59 26.01 9.22 $0.22\pm 0.11$ 0
2MASXJ12260880+2826008 186.5367 28.4336 186.5365 28.4348 8.18 64.95 9.12 $2.16\pm 1.15$ 1
PGC039730 184.9716 1.7734 184.974 1.7732 5.96 29.37 8.72 $0.39\pm 0.21$ 0
PGC039904 185.2971 17.6386 185.2966 17.6374 4.52 16.98 8.69 $0.18\pm 0.08$ 0
UGC07366 184.8695 17.2304 184.8706 17.2303 6.63 17.52 8.92 $0.15\pm 0.08$ 1
PGC560327 191.9905 -42.6486 191.9906 -42.6497 4.91 26.5 9.13 $0.24\pm 0.11$ 1
ESO324-008 200.2248 -39.2788 200.2239 -39.2774 4.72 27.12 8.99 $0.66\pm 0.19$ 0
HIZOAJ1343-65 205.8547 -65.2646 205.8548 -65.2651 5.37 43.61 8.94 $0.88\pm 0.37$ 1
LAMOSTJ134630.99+070428.6 206.6291 7.0746 206.6287 7.0747 2.41 110.43 9.17 $33.71\pm 5.98$ 1
ESO221-009 207.686 -48.9469 207.688 -48.9474 4.71 41.13 8.92 $1.2\pm 0.4$ 0
NGC5398 210.34 -33.0637 210.3406 -33.062 2.83 11.39 9.21 $0.23\pm 0.05$ 0
ESO271-018 212.5152 -46.2224 212.5127 -46.2235 5.01 35.29 8.15 $0.52\pm 0.24$ 0
PGC679199 212.6113 -33.2269 212.6111 -33.2282 5.01 50.8 9.4 $0.97\pm 0.47$ 1
PGC538542 218.2205 -44.3216 218.2217 -44.3212 4.41 15.51 8.65 $4.72\pm 0.32$ 1
ESO580-005 219.5665 -22.3245 219.5662 -22.3238 5.92 34.14 9.06 $0.59\pm 0.25$ 1
PGC812038 221.4301 -22.5793 221.4315 -22.5811 4.67 47.16 9.46 $13.42\pm 1.87$ 0
PGC053603 225.1221 -26.4499 225.1218 -26.4511 2.8 74.12 9.46 $7.49\pm 1.47$ 0
6dFJ1458243-373012 224.6013 -37.5031 $2$37.5028 1.74 104.8 9.33 $64.95\pm 7.9$ 0
PGC329372 309.7387 -63.771 309.74 -63.7704 5.39 21.35 8.44 $0.29\pm 0.13$ 1
2MASXJ21390659-4353001 324.$7$43.8832 324.7775 -43.8832 2.65 80.99 9.47 $22.01\pm 4.58$ 1
PGC3321235 332.5453 -56.0747 332.5466 -56.075 3.28 55.31 8.61 $5.99\pm 1.57$ 0
PGC3320749 330.1861 -56.6874 330.1838 -56.6863 4.92 63.14 8.81 $2.61\pm 1.16$ 0
NGC7657 351.698 -57.8058 351.6937 -57.8063 5.7 40.41 9.37 $0.74\pm 0.4$ 0
ESO240-012 354.7067 -51.8597 354.7074 -51.8597 5.15 23.69 9.41 $0.41\pm 0.16$ 1
ESO241-006 359.0627 -43.4275 359.0647 -43.4243 4.78 18.94 8.92 $0.29\pm 0.12$ 0
[…]
Appendix B X-ray archival data
B.1 Know ULXs
B.1.1 ESO 338-004
This galaxy is well known in the literature as the host of three ULXs, one of which might be an intermediate-mass black hole (IMBH) accretor (Oskinova et al., 2019).
B.1.2 SDSS J121855.02+142445.5
Another source is hosted in NGC 4254 and it is likewise thought to be powered by accretion onto an IMBH (Mezcua et al., 2013).
B.1.3 NGC 4038
One X-ray source is coincident with the nuclear region of the Antennae galaxy NGC 4038. This region has been imaged with both XMM-Newton and Chandra and its emission is attributed to the presence of several ULXs (Fabbiano et al., 2001).
B.1.4 PGC 014118
This galaxy location has been visited with XMM-Newton in February 2007. The X-ray emission appears to have originated from, at least, three different sources in the XMM-Newton image, spatially coincident with the location of a nuclear star cluster (NSC) (Fahrion et al., 2022). This suggests that the X-ray emission of this galaxy is caused by the blending of multiple ULXs, although the presence of an AGN cannot be ruled out.
B.2 Followed-up candidate AGN
B.2.1 NGC 3125
This source is a bright infrared LINER with a compact source of hard X-rays in its nucleus which is interpreted as the result of AGN activity (Dudik et al., 2005).
B.2.2 IC 549
This source’s location falls in the eFEDS footprint (Brunner et al., 2022) and was already indicated as an AGN candidate (Latimer et al., 2021). A follow-up observation by Chandra in 2022 further confirmed the AGN nature of this source.
B.2.3 PGC 039904
This source has been observed with Chandra and, given that its location lies outside the nucleus of its host galaxy, it is considered a ULX (Thygesen et al., 2023), although Bellovary et al. (2021) predicts that a fraction as large as 50% of AGN hosted in dwarf galaxies should be removed from its host nuclear region.
B.3 Bona fide AGN
B.3.1 PGC 053603
This galaxy was observed with XMM-Newton in February 2006 for $\approx 10$ ks. Its X-ray spectrum is well-modelled by an absorbed power law with photon index $\Gamma=2.3\pm 0.2$ and intrinsic absorption $N_{\textup{H}}=(1.5\pm 0.5)\times 10^{21}$ cm-2. The unabsorbed luminosity of the source in the $2-10$ keV band amounts to $(5.4\pm 0.9)\times 10^{40}\ \rm{erg\ s^{-1}}$.
B.3.2 PGC 538542
This galaxy was observed three times, in 2003, 2009 and 2014. We focus our analysis on this latest observation which is the longest and cleanest of the three, amounting to more than 90 ks. The X-ray spectrum of the nuclear region of this galaxy is well-modelled by a multi-component spectrum composed of a thermal plasma a power law and an emission Gaussian line at 5.9 keV. The thermal plasma temperature is $kT=1.17\pm 0.04$ keV, and the power law slope is $\Gamma=2.20\pm 0.08$. The unabsorbed luminosity of the sole power-law component in the $2-10$ keV band amount to $(3.7\pm 0.2)\times 10^{39}\ \rm{erg\ s^{-1}}$.
B.3.3 IC 2604
This galaxy has been observed with both XMM-Newton and Chandra and, although in both instruments its location falls more than $10^{\prime}$ from the observation pointing, the source is detected in both cases and the source location is roughly coincident with the nuclear region of its host galaxy. Its X-ray spectrum is well modelled by a power law with photon index $\Gamma=1.9\pm 0.1$ and the unabsorbed luminosity in the 2–10 keV band amounts to $L_{\textup{X}}=(5\pm 2)\times 10^{39}\ \rm{erg\ s^{-1}}$.
B.3.4 PGC 134076
This source has been serendipitously observed with Chandra. Its X-ray spectrum can be reproduced by a power law with photon index $\Gamma=1.5\pm 0.2$ and its X-ray luminosity in the 2–10 keV band is $L_{\textup{X}}=(4\pm 1)\times 10^{39}\ \rm{erg\ s^{-1}}$.
B.3.5 NGC 4016
The source location falls on the border of XMM-Newton detectors in a short and heavily affected by flaring observation in November 2020. Although the source is detected the handful of counts obtained does not allow for building a spectrum.
Appendix C Mock catalogs properties
The mock catalogs we built are based on the SAMs described in Ricarte & Natarajan (2018b), following the same approach of Chadayammuri et al. (2023). Figure 12 shows, for each SAM, the X-ray luminosity as a function of the stellar mass, and the cuts we imposed building our sample.
Figure 12: X-ray luminosity as a function of the stellar mass for each mock catalog we compared to our AGN sample. Shaded red dots indicate all the sources in our catalogs, and black dots are the ones that pass our selection criteria. The grey dotted lines show the threshold of our selection: no source with a luminosity below $10^{39}\ \rm{erg\ s^{-1}}$ or with a mass larger than $3\times 10^{9}~{}{\rm M}_{\odot}$ was considered.